A solid shaft of 10cm diameter transmits 74 kW at 150 rev/min. Calculate (a) the
torque on the shaft, (b) the maximum shear stress developed, (c) the angle of twist in a
length of 1.50m, and (d) the shear stress at a radius of 3cm. Take G = 8 MN/cm2
1
Expert's answer
2020-05-29T09:48:35-0400

a)


"P=\frac{2\pi NT}{60}\to 74000=\frac{2\pi (150)T}{60}"

"T=4711\ Nm"


b)


"T=\frac{\pi}{16}\tau d^3\to 4711=\frac{\pi}{16}\tau (0.1)^3"

"\tau=24\cdot10^6\frac{N}{m^2}"

c)


"\frac{T}{J}=\frac{G\theta }{l}\to \frac{T}{ \frac{\pi}{32} d^4}=\frac{G\theta }{l}"

"\frac{4711}{ \frac{\pi}{32} (0.1)^4}=\frac{8\cdot 10^{10}\theta }{1.5}"

"\theta=0.00899735\ rad=0.5155\degree"

d)


"T=\frac{\pi}{16}\tau' \frac{d^4-(2r)^4}{d}\to 4711=\frac{\pi}{16}\tau' \frac{0.1^4-(0.06)^4}{0.1}"

"\tau'=2.76\cdot10^7\frac{N}{m^2}"